OpenStudy (anonymous):

You buy a used car listed as $ 10,000 at a discount of 10 percent. However, you also pay 10 percent sales tax. Thus you end up writing a check for ? $

OpenStudy (anonymous):

car price after discount is: 10000*90/100= 9000 car+ 10% sales tax: 9000+ 10/100*9000=9000+900= $9900 i think this is the answer.